# Tweets to Citations: Unveiling the Impact of Social Media Influencers on AI Research Visibility

### v2

@@ 1 Introduction
Figure 1: The number of papers accepted to top AI/ML conferences (solid) and shared by influencers (dashed) from 2014-2023 Li (2023).
In the evolving landscape of artificial intelligence and machine learning (AI/ML), the exponential increase in conference papers, as depicted in Figure 1, alongside rapid technological advancements, has significantly transformed the dissemination of scholarly knowledge. A notable aspect of this transformation is the practice of online preprint sharing, with platforms like ArXiv becoming particularly prominent within the AI/ML community. This phenomenon allows for early access to research, often months before official publication, raising questions about the evolving relevance and reception of these papers in traditional academic forums. Most notably, how do people select papers to read in the online age?
This paper aims to explore the changing dynamics of academic discourse within the AI/ML community, emphasizing the constructive role of social media in addressing the challenges posed by the sheer volume of literature. We specifically focus on the case study of two influential $\mathbb{X}$ (formerly Twitter) users, AK (@_akhaliq) and Aran Komatsuzaki (@arankomatsuzaki), to understand how their social media activities aid in the curation and visibility of research. These influencers have emerged as pivotal figures in navigating the flood of information, akin to journalists in civic society, highlighting and contextualizing significant works for the community.
Acknowledging the vital function of these influencers as curators, we examine the impact of their endorsements on the citation counts of shared papers. However, we also underscore the importance of maintaining a balanced research ecosystem. An over-reliance on a select group of curators may inadvertently skew the research landscape, emphasizing certain topics or perspectives over others. Therefore, we advocate for a responsible approach to curation, encouraging influencers to maintain the journalistic standard that includes showcasing diverse research topics, authors, and institutions.
@@
$\bullet$ Thorough Analysis of Citations and Demographics. Our comprehensive analysis aims to determine if papers shared by these influencers receive higher citation counts compared to non-shared papers. We conduct statistical and causal inference analysis and find significant evidence for a direct effect. We then explore the geographic and gender distributions of influencer-shared papers and compare them to reference points from the community at large.
$\bullet$ Proposals for Future AI/ML Information Sharing. Finally, we propose that the academic community, particularly conference organizers, engage in a future discussion on evolving the conference system. This evolution is necessary to address the core challenge of managing an overwhelming number of submissions, ensuring that quality research is recognized and disseminated effectively. This approach will help ensure a more balanced and comprehensive understanding of research quality and relevance in these changing times.
@@ 2 Related Work
For over a decade now, social media platforms–namely $\mathbb{X}$–have been studied as an influential means of scholarly communication. Darling et al. (2013) discusses the potential role of Twitter in many stages of the ’life-cycle’ of academic authorship and publication. Eysenbach (2011) provides early evidence to support social media sharing as a predictor of higher citation count in medical publications. More recently, studies have shown a statistically significant relationship between Twitter presence and citation counts across fields, especially after the first tweet Peoples et al. (2016); Vaghjiani et al. (2021). Others suggest that the correlation between tweets and citations is insignificant through randomized trials Tonia et al. (2016); Branch et al. (2023).
@@ 3 Data Collection
Table 1: The number of unique papers tweeted by each influencer, the subset of papers with all desired attributes available through S2, and the papers finally included in our analysis after matching.
@@
Table 1: The number of unique papers tweeted by each influencer, the subset of papers with all desired attributes available through S2, and the papers finally included in our analysis after matching.
@@
We model our analysis on retrospective cohort studies, in which a treatment and control group with identical underlying covariates are compared to determine the average treatment effect. In our case, we assume that a paper’s citation count is most strongly influenced by elapsed time, quality, topic, and author prominence. While elapsed time is simple to measure, paper quality, topic are difficult to quantify. We use the publication venue and year as a proxy for quality and use a text embedding of the paper’s title and abstract to approximate the topic. We use first author total citation counts and h-index, and the maximum of all authors’ citation counts and h-indexes to adjust for their prominence.
@@
We find that roughly 90% of papers tweeted by either influencer have all attributes available. We express moderate caution about the minority of papers not included in our analysis, as they may be systematically different from papers we do include. However, we believe the large number of papers remaining in our analysis is sufficient to draw meaningful conclusions.
Table 3: The desired attributes for each paper, queried from S2.
@@
citationCount,
embedding.specter_v2,
@@
year, venue,
open_access, n_authors,
@@
Table 3: The desired attributes for each paper, queried from S2.
@@
Control Set. To build our Control Set, we first collect a large dataset of papers presented at the same venues and in the same years as those in the Target Set. Specifically, for every instance of a paper published in year $y$ at venue $v$, we query S2 for all papers published in year $y$ at venue $v$.
We follow the recommendations in King and Nielsen (2019) by performing exact matching on our categorical (year, venue, open_access [^1] Due to the nature of influencer sharing, all sampled papers are open access.) and binned variables (n_authors, h-index, author cites), and Euclidean distance [^2] Note that Euclidean distance is equivalent to cosine similarity for normalized vectors. matching on our continuous variables (topic_embedding). Optimal matching can be reduced to the linear sum assignment problem, for which many efficient algorithms exist Crouse (2016). We use the implementation available in SciPy Virtanen et al. (2020).
Using this methodology, we match each paper in the Target Set to a paper in the Control Set with respect to the gathered covariates (Table 4). We exclude any paper for which we cannot find an exact match on categorical variables. For matching on topic embeddings, we use SPECTER2 Singh et al. (2022). Matches between papers were strong, and the distribution of cosine similarities as well as example matched pairs are detailed in Appendix A.
@@
We plot the mean review score of each treatment paper against the mean score of its paired control (Figure 2). For conferences that do not use numbered review scores, we assign numbers based on those of other conferences (7: Accept, 5: Borderline Accept, 3: Borderline Reject, 2: Reject).
@@ 4 Analysis
We compare our paired target and control sets, as described in Section 3. We find that papers tweeted by AK have a median citation count of 24 (95% CI: 23, 25) versus 14 (95% CI: 13, 15) in the control group, and papers tweeted by Komatsuzaki have a median citation count of 31 (95% CI: 27, 34) versus 12 (95 % CI: 10.5, 13.5). Visually, we can see that both experimental set distributions are skewed toward higher citation counts when compared to their corresponding control sets (Figure 3). In the violin plots (Figures 2(c) and 2(d)), the three quartiles and max values are all higher in both of the shared paper distributions compared to the controls. In the 2-Sample Q-Q plots (Figure 4), we can see that the normalized quantiles are consistently higher for the test distributions.
@@
Overall, the correlation between influencer tweets and citation count–and not review scores–points to a shift in how the community finds and reads papers. While traditionally, top conference acceptance (i.e. review score) has been a primary indicator of future citation count Lee (2018), we have shown that the sharing practices of far-reaching influencers are now a significant indicator of future research impact through citations.
(a) @_akhaliq Dataset
(b) @arankomatsuzaki Dataset
@@
We aim to estimate the causal effect of the treatment $A$, indicating that the paper was shared by influencers [^3] To increase sample size and simplify our analysis, we combine our test sets for this portion of the analysis., on the outcome $Y$, if the paper is "highly-cited" or "less-cited." Besides $A$, there are a number of other factors, known as confounders, that can affect $Y$. We divide these confounders into two sets: observed $C$–which are the same used in our matching (Section 3)–and unobserved $U$–which are difficult to quantify or measure (e.g. novelty, contribution, hype).
To debias the effects of unobserved confounders $U$, we employ the negative outcome control (NOC) framework Card and Krueger (1993); Lipsitch et al. (2010). By finding a negative control outcome $N$ that shares the same confounders as $Y$, but is not causally affected by $A$, we can attempt to correct for the bias introduced by $U$. To this end, we select a paper’s average review score for $N$, which we have already shown is not significantly correlated to $A$ (Figure 2) and which we believe share many of the same confounders (e.g., quality, contribution, author experience). However, another NCO can be substituted if deemed more suitable.
@@
Finally, we estimate the causal effect of influencer sharing on paper citations with the average treatment effect of the treated. Using the difference-in-difference assumption with NOC, we can write our estimate Elazar et al. (2023):
@@
Figure 5: Forest plot of ATET and confidence intervals approximated with negative outcome control (NOC). Larger positive values indicate a stronger positive causal effect of the treatment (influencer sharing) on the outcome (a paper being "highly-cited"). To ensure robust results, we vary the quantile threshold for "highly-cited" and "highly-scored" papers; "Unadj." values (red squares) show the effect estimate before applying NOC. These results, where no confidence intervals contain 0%, indicate a significant positive causal effect of influencer sharing on paper citations.
@@
We must note, though, that using solely self-reported affiliations can have an inherent bias toward the United States. For example, many researchers affiliated with U.S.-based organizations are assigned to the United States despite working out of another area. Additionally, we must note the prominence of the "Unknown" category in the data of both influencers, where affiliations were not found.
@@
Figure 7: AI repository publications by geographic area and publication year, 2010–2021. Data extracted from Figure 1.1.19 of the Stanford HAI 2023 AI Index Report Maslej et al. (2023).
@@
Gender Distributions. Beyond geographic diversity, gender diversity is crucial in Computer Science and Engineering, fields historically dominated by men. We extract author names and affiliations as described above. In this section, we filter only the first authors of each paper. For gender classification, we used the AMiner Scholar Gender Prediction API, which categorizes authors as "male," "female," or "UNKNOWN" based on name and affiliation–if available. The API uses a majority vote of the results from three sources: google image search and facial recognition, the Facebook Generated Names List Tang et al. (2011), and WebGP Gu et al. (2016).
To ground our view of the overall gender distribution in the field, we reference the Taulbee survey’s reported gender distribution of US Ph.D. awardees and faculty in CS and related fields from 2021-2022 Zweben and Bizot (2023). To match the classifications we have available, we will consider the binary reported genders from the survey.
@@ 5 Discussion
However, we advise caution against excessive dependence on a limited set of information sources. Research inherently involves bottom-up exploration of a wide array of topics. Focusing on a handful of individuals’ highlighted papers necessarily offers a narrow view of the research landscape. We encourage the community to keep the online academic space competitive, allowing for a diversity of highly visible ideas. This goal can be achieved through active participation in an open, community-driven curation process, enhancing the variety of prominent ideas. Additionally, we urge influencers to share a diverse array of ideas. This includes showcasing various techniques and subtopics within their areas of focus, thereby exposing community members to new approaches and concepts.
Enhancing Equity. Unquestionably, we would rather have more voices than fewer in the machine learning community. The trend towards influencer-led dissemination in the ML community offers new avenues to enhance equity. Our geographic analysis indicates a bias towards U.S.-centered research in influencer-shared papers. Despite reflecting the current U.S. prominence in AI/ML repository publications, it also reveals the potential to include research from a broader range of regions, including Europe, Asia, Africa, and Latin America. Similarly, Machine Learning sees a dramatic disparity in male versus female authors. Additionally, our gender distribution analysis, though not indicating a bias in influencer-shared content, highlights the general gender disparity within ML. This disparity presents an opportunity to foster gender diversity in the field.
While we specifically examine geography and gender, these principles apply broadly: proactive efforts in the online domain can significantly enhance equity. We encourage influencers to actively promote visibility for diverse groups globally. We emphasize that this does not necessarily indicate bias among current influencers but serves as a call for vigilance against potential biases as the AI/ML domain expands. This connects back to our recommendation for fostering a diverse online space where everyone can contribute to broadening the spectrum of shared ideas.
@@ 6 Conclusion
This study delves into the influence of social media influencers on the dissemination and recognition of academic work in AI/ML, specifically examining the impact of AK and Aran Komatsuzaki on platforms like $\mathbb{X}$ and Hugging Face. Our analysis demonstrates that papers endorsed by these influencers receive significantly more citations than those that are not, underscoring the role of influencers in not only amplifying the reach of specific research but also in shaping its visibility in the field. This influence extends beyond merely sharing higher-quality papers, highlighting their ability to effectively contextualize and promote substantial findings within the community.
The study further discusses the dual role of influencers as both catalysts for visibility and curators of content, emphasizing the need for a balance in their influence to ensure a diversity of perspectives in the AI/ML research landscape. The implications of this influence are profound, suggesting a need for the academic community to reassess traditional methods of paper selection and review. Our findings encourage a dialogue among conference organizers and academic institutions to evolve the peer-review process and adapt to these changing norms. Additionally, the study opens avenues for future research into the generalizability of these trends in other scientific fields and the mechanisms behind the influence of social media on academic recognition.
@@ Appendix A Target-Control Matching
Score: 0.09
Target Title: Unifying Vision, Text, and Layout for Universal Document Processing
@@
Score: 0.09
Target Title: reStructured Pre-training
@@
Score: 0.12
Target Title: Quantifying Memorization Across Neural Language Models
@@
Score: 0.07
Target Title: Do Language Models Know When They’re Hallucinating References?
Target Abstract: State-of-the-art language models (LMs) are famous for"hallucinating"references. These fabricated article and book titles lead to harms, obstacles to their use, …
@@
Score: 0.10
Target Title: Singularity: Planet-Scale, Preemptive and Elastic Scheduling of AI Workloads
@@
Score: 0.09
Target Title: A Large-Scale Study on Unsupervised Spatiotemporal Representation Learning

### v3

# Position: AI/ML Influencers Have a Place in the Academic Process

@@ 1 Introduction
Figure 1: The number of papers accepted to top AI/ML conferences (solid) and shared by influencers (dashed) from 2014-2023 (Li, 2023).
In the evolving landscape of artificial intelligence and machine learning (AI/ML), the exponential increase in conference papers, as depicted in Figure 1, alongside rapid technological advancements, has significantly transformed the dissemination of scholarly knowledge. A notable aspect of this transformation is the practice of online preprint sharing, with platforms like ArXiv becoming particularly prominent within the AI/ML community. This phenomenon allows for early access to research, often months before official publication, raising questions about these papers’ evolving relevance and reception in traditional academic forums. Most notably, how do people select papers to read in the online age?
This paper aims to explore the changing dynamics of academic discourse within the AI/ML community, emphasizing the constructive role of social media in addressing the challenges posed by the sheer volume of literature. We study the class of $\mathbb{X}$ (formerly Twitter) influencers. Specifically, we require that influencers (1) share freshly posted preprints from ArXiv, (2) regularly post about papers, (3) do not focus on self-promotion, and (4) have a large audience. AK (@_akhaliq) and Aran Komatsuzaki (@arankomatsuzaki) are the only influencers found to satisfy our requirements.
Acknowledging the vital function of these influencers as curators, we examine the impact of their endorsements on the citation counts of shared papers. However, we also underscore the importance of maintaining a balanced research ecosystem. An over-reliance on a select group of curators may inadvertently skew the research landscape, emphasizing certain topics or perspectives over others. Therefore, we advocate for a responsible approach to curation, encouraging influencers to maintain journalistic standards, showcasing diverse research topics, authors, and institutions.
@@
$\bullet$ Thorough Analysis of Citations and Demographics. Our comprehensive analysis aims to determine if papers these influencers share receive higher citation counts than non-shared papers. We conduct statistical and causal inference analysis and find significant evidence for a direct effect. We then explore the geographic and gender distributions of influencer-shared papers and compare them to reference points from the community.
$\bullet$ Proposals for Future AI/ML Information Sharing. Finally, we propose that the academic community engage in a future discussion on evolving the conference system. We want to collaborate with program chairs at major ML conferences to find solutions to the AI/ML research cycle outpacing current conference cycles, making many works outdated by the time of conferences. Second, we propose a future workshop to address the issue by informing the AI/ML community and gathering solutions from community members. Third, we call for a panel of influencers, industry, and academics to discuss how the evolving landscape of AI/ML research can be improved.
@@ 2 Related Work
For over a decade now, social media platforms–namely $\mathbb{X}$–have been studied as an influential means of scholarly communication. Darling et al. (2013) discusses the potential role of Twitter in many stages of the ’life-cycle’ of academic authorship and publication. Eysenbach (2011) provides early evidence to support social media sharing as a predictor of higher citation count in medical publications. More recently, studies have shown a statistically significant relationship between Twitter presence and citation counts across fields, especially after the first tweet (Peoples et al., 2016; Vaghjiani et al., 2021). Others suggest that the correlation between tweets and citations is insignificant through randomized trials (Tonia et al., 2016; Branch et al., 2023).
@@ 3 Data Collection
Table 1: The number of unique papers tweeted by each influencer, the subset of papers with all desired attributes available through S2 (Table 4), and the papers finally included in our analysis after matching.
We model our analysis on retrospective cohort studies, in which a treatment and control group with identical underlying covariates are compared to determine the average treatment effect. In our case, we assume that a paper’s citation count is most strongly influenced by elapsed time, quality, topic, and author prominence. While elapsed time is simple to measure, paper quality and topic are difficult to quantify. We use the publication venue and year as a proxy for quality and use a text embedding of the paper’s title and abstract to approximate the topic. We use author citations and h-indices to adjust for their prominence.
@@
We find that roughly 90% of papers tweeted by either influencer have all attributes available. While we express moderate caution about the excluded minority of papers, as they may be systematically different from those included, we believe the many papers remaining in our analysis are sufficient to draw meaningful conclusions.
Additionally, we find that half of the papers are tweeted within 24 hours of their initial ArXiV release, with 95% of tweets occurring within a week. This verifies that our influencers do indeed tweet freshly posted papers. We also find significant overlap in sharing, with 65% of Komatsuzaki-shared papers also tweeted by AK.
@@
citationCount, embedding.specter_v2,
@@
Table 3: The desired attributes for each paper, queried from S2.
year, venue, open_access, n_authors,
@@
Control Set. To build our Control Set, we first collect a corpus of papers presented at the same venues and in the same years as those in the Target Set. Specifically, for every instance of a paper published in year $y$ at venue $v$, we query S2 for all papers published in year $y$ at venue $v$.
We follow the recommendations in (King & Nielsen, 2019) by performing exact matching on our categorical (year, venue, open_access [^1] Due to the nature of influencer sharing, all sampled papers are open access.) and binned variables (n_authors, h-index, author cites), and Euclidean distance [^2] Note that Euclidean distance is equivalent to cosine similarity for normalized vectors. matching on our continuous variables (topic_embedding). Optimal matching can be reduced to the linear sum assignment problem, for which many efficient algorithms exist (Crouse, 2016). We use the implementation available in SciPy (Virtanen et al., 2020).
Using this methodology, we match each paper in the Target Set to a paper in the Control Set with respect to the gathered covariates (Table 4). We exclude any paper for which we cannot find an exact match on categorical variables. For matching on topic embeddings, we use SPECTER2 (Singh et al., 2022). Matches between papers were strong, and the distribution of cosine similarities and example matched pairs are detailed in Appendix A.
@@
We plot each treatment paper’s mean review score against its paired control’s mean score (Figure 2). For conferences that do not use numbered review scores, we assign numbers based on those of other conferences (7: Accept, 5: Borderline Accept, 3: Borderline Reject, 2: Reject).
@@ 4 Analysis
We compare our paired target and control sets, as described in Section 3. We find that papers tweeted by AK have a median citation count of 24 (95% CI: 23, 25) versus 14 (95% CI: 13, 15) in the control group, and papers tweeted by Komatsuzaki have a median citation count of 31 (95% CI: 27, 34) versus 12 (95 % CI: 10.5, 13.5). Visually, we can see that both experimental set distributions are skewed toward higher citation counts when compared to their corresponding control sets (Figure 3). In the violin plots (Figures 3(c) and 3(d)), the three quartiles and max values are all higher in both of the shared paper distributions compared to the controls. In the 2-Sample Q-Q plots (Figure 4), we can see that the normalized quantiles are consistently higher for the test distributions.
@@
Overall, the correlation between influencer tweets and citation count–and not review scores–points to a shift in how the community finds and reads papers. While traditionally, top conference acceptance (i.e. review score) has been a primary indicator of future citation count (Lee, 2018), we have shown that the sharing practices of far-reaching influencers are now a significant indicator of future research impact through citations.
Figure 5: Forest plot of ATET and confidence intervals approximated with negative outcome control (NOC). Larger positive values indicate a stronger positive causal effect of the treatment (influencer sharing) on the outcome (a paper being ”highly-cited”). To ensure robust results, we vary the quantile threshold for ”highly-cited” and ”highly-scored” papers; ”Unadj.” values (red squares) show the effect estimate before applying NOC. These results, where no confidence intervals contain 0%, indicate a significant positive causal effect of influencer sharing on paper citations.
@@
We aim to estimate the causal effect of the treatment $A$, indicating that the paper was shared by influencers [^3] To increase sample size and simplify our analysis, we combine our test sets for this portion of the analysis., on the outcome $Y$, if the paper is ”highly-cited” or ”less-cited.” Besides $A$, there are a number of other factors, known as confounders, that can affect $Y$. We divide these confounders into two sets: observed $C$–which are the same used in our matching (Section 3)–and unobserved $U$–which are difficult to quantify or measure (e.g. novelty, contribution, hype).
To debias the effects of unobserved confounders $U$, we employ the negative outcome control (NOC) framework (Card & Krueger, 1993; Lipsitch et al., 2010). By finding a negative control outcome $N$ that shares the same confounders as $Y$, but is not causally affected by $A$, we can attempt to correct for the bias introduced by $U$. To this end, we select a paper’s average review score for $N$, which we have already shown is not significantly correlated to $A$ (Figure 2) and which we believe share many of the same confounders (e.g., quality, contribution, author experience). However, another NCO can be substituted if deemed more suitable.
@@
Finally, we estimate the causal effect of influencer sharing on paper citations with the average treatment effect of the treated. Using the difference-in-difference assumption with NOC, we can write our estimate (Elazar et al., 2023):
@@
Figure 7: AI repository publications by geographic area and publication year, 2010–2021. Data extracted from Figure 1.1.19 of the Stanford HAI 2023 AI Index Report (Maslej et al., 2023).
@@
We must note, though, that using solely self-reported affiliations can have an inherent bias toward the United States. For example, many researchers affiliated with U.S.-based organizations are assigned to the United States despite working out of another area. Additionally, we must note the prominence of the ”Unknown” category in the data of both influencers, where affiliations were not found.
@@
Gender Distributions. Beyond geographic diversity, gender diversity is crucial in Computer Science and Engineering, fields historically dominated by men. We extract author names and affiliations as described above. In this section, we filter only the first authors of each paper. For gender classification, we used the AMiner Scholar Gender Prediction API, which categorizes authors as ”male,” ”female,” or ”UNKNOWN” based on name and affiliation–if available. The API uses a majority vote of the results from three sources: google image search and facial recognition, the Facebook Generated Names List (Tang et al., 2011), and WebGP (Gu et al., 2016).
To ground our view of the overall gender distribution in the field, we reference the Taulbee survey’s reported gender distribution of US Ph.D. awardees and faculty in CS and related fields from 2021-2022 (Zweben & Bizot, 2023). To match the classifications we have available, we will consider the binary reported genders from the survey.
@@ 5 Discussion
Yet, reliance on a select few narrows the full spectrum of research. We propose maintaining a competitive, open academic space with diverse, highly visible ideas, advocating for influencers to diversify showcased techniques and concepts. This encourages a broad exploration of the ML landscape, mitigating the risk of a homogeneous research narrative.
Enhancing Equity. The move towards influencer-led dissemination opens paths to greater equity in the ML community. Our findings—pointing out the geographic and gender disparities in shared research—underscore the potential for a more inclusive academic dialogue. We clarify that our analysis does not suggest influencers currently exhibit bias; rather, it highlights an opportunity to mitigate existing inequities through proactive online engagement. By promoting diverse global perspectives and addressing gender disparities, influencers can foster a more equitable field. This effort aligns with our commitment to enhance equity without overemphasizing certain demographics, reflecting the concentration of ML research.
To stimulate community reflection and action, we recommend collaborating with conference chairs to address the pace of research, organizing workshops to explore dissemination improvements, and convening panels of influencers, industry, and academia to discuss enhancements in AI/ML research visibility. These steps aim to refine the research dissemination process and uplift the ICML community’s engagement with emerging and diverse ideas.
@@ 6 Conclusion
Our study delves into the influence of social media influencers on the dissemination and recognition of academic work in AI/ML, specifically examining the impact of AK and Aran Komatsuzaki on platforms like $\mathbb{X}$ and Hugging Face. Our analysis demonstrates that papers endorsed by these influencers receive significantly more citations than those that are not, underscoring the role of influencers in not only amplifying the reach of specific research but also in shaping its visibility in the field. This influence extends beyond merely sharing higher-quality papers, highlighting their ability to promote substantial findings within the community.
We further discuss the dual role of influencers as both catalysts for visibility and curators of content, emphasizing the need for a balance in their influence to ensure a diversity of perspectives in the AI/ML research landscape. The implications of this influence are profound, suggesting a need for the academic community to reassess traditional methods of paper selection and review. Our findings encourage a dialogue among conference organizers and academic institutions to evolve the peer-review process and adapt to these changing norms. Additionally, we encourage future research into these trends in other scientific fields and the mechanisms behind the influence of social media on academic recognition.
@@
Acknowledgements
This work was supported in part by the National Science Foundation Research Experience for Undergraduates under CAREER Award Grant No. 2048122.
@@ Appendix A Target-Control Matching
Score: 0.09 Target Title: Unifying Vision, Text, and Layout for Universal Document Processing
@@
Score: 0.09 Target Title: reStructured Pre-training
@@
Score: 0.12 Target Title: Quantifying Memorization Across Neural Language Models
@@
Score: 0.07 Target Title: Do Language Models Know When They’re Hallucinating References?
Target Abstract: State-of-the-art language models (LMs) are famous for”hallucinating”references. These fabricated article and book titles lead to harms, obstacles to their use, …
@@
Score: 0.10 Target Title: Singularity: Planet-Scale, Preemptive and Elastic Scheduling of AI Workloads
@@
Score: 0.09 Target Title: A Large-Scale Study on Unsupervised Spatiotemporal Representation Learning